OpenStudy (anonymous):

3/4(8+4x) - 1/3(6x+3)= Solve using common denominators.

OpenStudy (anonymous):

So, just to make sure, this is the problem you're looking for, right?: \[\frac{ 3 }{ 4 } (8+4x) - \frac{ 1 }{ 3 }(6x+3)=0\]

OpenStudy (anonymous):

no sorry it equals 9 :P

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }(8+4x) - \frac{ 1 }{ 3 }(6x+3)= 9\]

OpenStudy (anonymous):

Okay great! :) So, to start off, you would want to distribute the '3' in the 3/4 to (8+4x). So, that would be: \[\frac{ 24+12x }{ 4 }\] Then, you would do the same to the other fraction- it would be: \[\frac{ 6x+3 }{ 3 }\] Now, you want to get common denominators. Do you think you've got it from here? If not, let me know. :)

OpenStudy (anonymous):

Oh i see. I think I got it. But wait a sec, let me check if i can get the answer. ;D

OpenStudy (anonymous):

Okay, just let me know what you get and I'll tell you if it's right!

OpenStudy (anonymous):

\[\frac{ 48+12x }{ 12 }\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

so its : \[4+x\]

OpenStudy (anonymous):

wait i fogot there is a 9. Sorry

OpenStudy (anonymous):

forgot*

OpenStudy (anonymous):

so is the answer x=5?

OpenStudy (anonymous):

Well, not quite- You want to multiply the top and bottom of 24+12x/4 by 3, so it would be 72+36x/12. And you'd want to multiply the top and bottom of 6x+3/3 by 4, so it would be 24x+12/12 ... Normally! But you can't forget that the two are being subtracted, so you have to make the second one -24x-12/12. Then, just put the tops together, simplify, and you're almost there! \[\frac{ 72+36x-24x-12 }{ 12 }\] Simplify further... \[\frac{ 60+12x }{ 12 }\] And don't forget that you can take '12' out of '60', so this simplifies up beautifully to: \[\frac{ 12(5+x) }{ 12 }\] And from that to be: 5+x=9 -5 -5 x=4.

OpenStudy (anonymous):

oh. Yes i see now. Thank You! :D

OpenStudy (anonymous):

i forgot to change the 12 to a negative

OpenStudy (anonymous):

No problem! :D